Pantheon Operations
All Systems Operational
Customer Sites ? Operational
Dashboard ? Operational
Workflow Operations ? Operational
Spinup Operations ? Operational
Support Tickets ? Operational
Support Chat ? Operational
Global CDN ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
System Metrics Month Week Day
Customer Site Availability
Fetching
Dashboard Response Time
Fetching
Past Incidents
May 22, 2017

No incidents reported today.

May 21, 2017

No incidents reported.

May 20, 2017

No incidents reported.

May 19, 2017
Resolved - All clear. Our upstream provider reports issuing certificates is now fully operational, which our monitoring confirms.
May 19, 16:24 PDT
Monitoring - An upstream provider is experiencing issues, domains on Pantheon's HTTPS certificates may take longer than normal to add or remove.
May 19, 08:46 PDT
Identified - An upstream provider is having intermittent problems and they are working on fixing them.
May 19, 07:48 PDT
Monitoring - A fix has been implemented and we are monitoring the results.
May 19, 06:40 PDT
Identified - The issue has been identified and a fix is being implemented.
May 19, 05:51 PDT
Monitoring - A fix has been implemented and we are monitoring the results.
May 19, 04:38 PDT
Identified - An upstream provider is experiencing issues, domains cannot be added or removed from Pantheon's HTTPS certificates at this time.
May 19, 02:39 PDT
Resolved - Solr service is back to normal. Incident is resolved. Thank you for your patience!
May 19, 12:53 PDT
Monitoring - Solr service is stable and we will continue to monitor performance.
May 19, 12:01 PDT
Investigating - We are investigating degraded Apache Solr performance. Thank you for your patience.
May 19, 11:05 PDT
May 18, 2017
Resolved - Solr service is back to normal. Incident is resolved. Thank you for your patience!
May 18, 15:42 PDT
Monitoring - Solr service is stable and we will continue to monitor performance.
May 18, 15:05 PDT
Investigating - We are investigating degraded Apache Solr performance.
May 18, 14:14 PDT
Resolved - Connectivity issues have cleared. Incident resolved. Thanks again for your patience.
May 18, 11:36 PDT
Monitoring - The connectivity issues appear to have been resolved and we are continuing to monitor.
May 18, 09:57 PDT
Update - We are continuing to investigate this incident.
May 18, 08:33 PDT
Investigating - We are investigating connectivity issues and degraded performance around IPv6 addresses for DNS.
May 18, 06:49 PDT
May 17, 2017
Resolved - This incident has been resolved.
May 17, 13:42 PDT
Monitoring - We have remediated the issue and are now monitoring the affected services.
May 17, 12:56 PDT
Identified - We have identified an issue with an index server that is causing partial search outage for a portion of customer sites.
May 17, 12:15 PDT
May 16, 2017

No incidents reported.

May 15, 2017
Resolved - This incident has been resolved.
May 15, 15:10 PDT
Monitoring - We have implemented a fix and are monitoring the stability at this time. Thank you for your patience.
May 15, 13:48 PDT
Investigating - We are currently investigating this issue.
May 15, 13:15 PDT
May 14, 2017

No incidents reported.

May 13, 2017

No incidents reported.

May 12, 2017

No incidents reported.

May 11, 2017

No incidents reported.

May 10, 2017

No incidents reported.

May 9, 2017

No incidents reported.

May 8, 2017

No incidents reported.